### v1

# Semantic Superiority vs. Forensic Efficiency: A Comparative Analysis of Deep Learning and Psycholinguistics for Business Email Compromise Detection

Abstract
Business Email Compromise (BEC) is a sophisticated social engineering threat that manipulates organizational hierarchies and exploits psychological vulnerabilities like trust and authority, leading to significant financial and reputational damage for enterprises. According to the 2024 FBI Internet Crime Report, BEC accounts for over \$2.9 billion in annual adjusted losses, presenting a significant economic asymmetry: the cost of a False Negative (fraud loss) exceeds the cost of a False Positive (manual review) by orders of magnitude (approximately 1 to 5,480).
This paper examines two distinct detection paradigms for BEC: the Forensic Psycholinguistic Stream, which utilizes CatBoost to analyze psycholinguistic cues with high interpretability and low latency, and the Semantic Stream, which employs DistilBERT for deep learning-based contextual language understanding, offering superior accuracy at the cost of higher computational resources. We evaluated DistilBERT on an adversarially poisoned dataset (N = 7,990) generated via our Black Hole protocol. It was benchmarked on Tesla T4 GPU infrastructure, achieving superior detection (AUC = 1.0000, F1 = 0.9981) with acceptable real-time latency (7.403 milliseconds). CatBoost achieves competitive detection (AUC = 0.9905, F1 = 0.9486) at 8.4x lower latency (0.885 milliseconds), consuming negligible computational resources. For organizations with GPU infrastructure, DistilBERT offers superior accuracy, making it an ideal choice. On the other hand, CatBoost is preferable for edge deployments or cost-sensitive environments due to its comparable security and lower operational costs, providing a viable alternative. Both approaches demonstrate a return on investment exceeding 99.96% when optimized through cost-sensitive learning, as evidenced by their ability to significantly reduce false negatives and associated financial losses. This high return is achieved by minimizing the financial impact of undetected fraudulent emails, which are substantially more costly than false positives.
I Introduction
Business Email Compromise (BEC) represents a shift in cybercrime tactics. Unlike ransomware or malware, which exploit technical vulnerabilities (e.g., buffer overflows, zero-day exploits), BEC exploits human cognition. Attackers use social engineering to manipulate trust, expanding the attack surface from “CEO Fraud” to complex “Vendor Email Compromise” (VEC) and “Payroll Diversion,” compromising legitimate accounts to send fraudulent invoices.
I-A Background and Motivation
These attacks have become more sophisticated with Generative AI. Attackers can generate mathematically unique but semantically identical pretexts at scale using Large Language Models (LLMs), rendering hash-based signature detection obsolete [4]. Furthermore, Unicode homoglyphs (e.g., substituting Latin ‘a’ with Cyrillic ‘a’) create attacks that appear identical to humans but break the tokenization logic of standard NLP models [5].
As BEC tactics become increasingly sophisticated, defenders are confronted with a pivotal decision: to employ Transformer-based models like BERT, which offer superior contextual understanding but require significant computational resources, or to choose lightweight feature-engineered models that, while efficient, provide less contextual awareness. This decision is crucial as it impacts the balance between detection accuracy and resource allocation.
I-B Problem Statement
Current academic literature views BEC detection as a text classification problem focused on metrics like Accuracy or F1-Score, overlooking two operational realities:
  1. 1.
    Economic Asymmetry: The cost of a False Positive ($C_{FP}$) is approximately \$25. The cost of a False Negative ($C_{FN}$) is approximately \$137,000 on average [1]. Standard loss functions fail to capture this 1:5,480 risk ratio.
  2. 2.
    Infrastructure Constraints: GPU deployment improves Transformer inference latency. However, sustained GPU access is often limited. A cost-benefit analysis is necessary; DistilBERT demands significant GPU resources, whereas CatBoost operates with minimal memory.
I-C Paper Organization
The paper is structured to guide the reader through a comprehensive exploration, beginning with related work and the threat model to establish the foundation, followed by the theoretical framework and methodology to detail the approach. It then dives into the experimental setup and results to present findings, examines forensic failure modes to identify potential weaknesses, discusses economic optimization to highlight cost-effectiveness, and concludes with key findings to summarize the study’s contributions.
II Related Work
II-A NLP in Phishing Detection
Early work focused on structural metadata and keyword density [6]. While effective against “spray-and-pray” phishing, these methods fail against targeted BEC where the attacker mimics standard business protocols. The advent of Transformers (BERT, RoBERTa) shifted focus to contextual embeddings. Lee et al. [7] demonstrated high accuracy with BERT on general phishing, but their work did not account for the “clean” nature of BEC, which lacks malicious URLs or attachments. Sanh et al. [8] introduced DistilBERT, offering a viable latency-performance tradeoff for real-time email scanning, which serves as the backbone for our semantic stream.
II-B Adversarial Text Attacks
NLP models face challenges from character-level changes like Unicode obfuscation and homoglyph attacks, which can lead to errors in tokenization and misclassification. Ebrahimi et al. [9] demonstrated “HotFlip,” a method for generating adversarial examples by swapping characters. In BEC, attackers use Unicode obfuscation (homoglyphs) to evade tokenizers [11]. Standard BERT models treat “Bank” and “Bank” (Cyrillic ‘a’) as entirely unrelated tokens.
II-C Cost-Sensitive Learning
Elkan [12] established the foundations of cost-sensitive learning, arguing that class probability thresholds should shift based on misclassification costs. However, few papers address BEC. Our work integrates a dynamic Value-at-Risk (VaR) model into the decision logic, similar to credit fraud detection in high-frequency trading [13].
III Threat Model
III-A Security Assumptions
While standard authentication protocols (SPF, DKIM, DMARC) make spoofing the CEO’s exact domain challenging for attackers, they may still use look-alike domains or compromised vendor accounts that pass SPF/DKIM.
III-B Risk Scenarios (MITRE ATT&CK)
We focus on three primary attack vectors in the MITRE ATT&CK framework:
  • T1598.002 (Spearphishing via Service): Impersonating an executive to demand urgent wire transfers.
  • T1566.002 (Spearphishing Link): Embedding QR codes (Quishing) to bypass text scanners.
  • T1036.005 (Masquerading): Using homoglyphs to mimic legitimate keywords (e.g., “Payment”).
IV Theoretical Framework
IV-A Psycholinguistics: Grice’s Maxims & The “Smiling Assassin”
Legitimate business communication typically adheres to Grice’s Maxim of Quality (be truthful). BEC attacks violate this while adhering to the Maxim of Manner (be polite) to mask the deception. We quantify this “Polite Coercion” using the Smiling Assassin Score ($\psi$).
Let $S_{pos}(x)\in[0,1]$ be the normalized positive sentiment and $U_{\text{freq}}(x)$ be the urgency token density.
$\psi(x)=\sigma\left(\alpha\cdot S_{pos}(x)\cdot\ln(1+\beta\cdot U_{\text{freq}}(x))\right)$ (1)
Where $\sigma$ is the logistic sigmoid function.
Sensitivity Analysis: The derivative $\frac{\partial\psi}{\partial U}$ is maximized when $S_{pos}\to 1$. This models the attacker’s attempt to lower defenses via politeness before introducing urgency (“grooming”).
IV-B Financial Loss Function ($L_{fin}$)
We replace standard Cross-Entropy with a financial loss function $L_{fin}$ to optimize decision thresholds.
$E[L_{fin}]=\sum_{i=1}^{N}\Big(\mathbb{I}_{FN}\cdot V_{i}+\mathbb{I}_{FP}\cdot C_{inv}+\mathbb{I}_{G}\cdot C_{rev}\Big)$ (2)
Where:
  • $V_{i}$: Estimated transaction value (default \$137k based on IC3 data [1]).
  • $C_{inv}$: Investigation cost for a blocked legitimate email ($\approx\mathdollar 25$).
  • $C_{rev}$: Human review cost for ambiguous emails ($\approx\mathdollar 25$).
V Methodology
Figure 1: Comparative Architecture. Stream A (Forensic/Green) extracts psycholinguistic features for a Gradient Boosting classifier. Stream B (Semantic/Blue) uses a hardened Tokenizer and Transformer to extract deep embeddings. GPU acceleration reduces DistilBERT latency to real-time acceptable ranges.
V-A Data Engineering: The “Black Hole” Protocol
We constructed a hybrid dataset ($N=7,990$) combining:
  1. 1.
    Legitimate: Enron Corpus [15], processed via V7 Anonymization to remove PII and mitigate temporal drift.
  2. 2.
    Adversarial: GPT-4 generated 4,211 synthetic BEC emails based on 30 taxonomies.
  3. 3.
    Poison Injection: 30% of fraud samples were injected with Homoglyphs (Cyrillic ‘a’) and Zero-Width Spaces to simulate T1036 Masquerading.
V-B Adversarial Hardening Algorithm
To ensure the Semantic Stream is robust, we implemented a pre-processing normalization algorithm (Algorithm 1). This prevents the tokenizer from splitting “Bank” into nonsense tokens.
Algorithm 1 Homoglyph Normalization Strategy
0: Raw Email Text $T$, Homoglyph Map $M$
0: Normalized Text $T_{norm}$
[…]
2:for each character $c$ in $T$ do
3:  if $c$ is in $M$.keys() then
4:   $c_{safe}\leftarrow M[c]$
5:   Append $c_{safe}$ to $T_{norm}$
6:  else
7:   if $c$ is not Invisible (e.g., U+200B) then
8:    Append $c$ to $T_{norm}$
9:   end if
10:  end if
11:end for
12:return $T_{norm}$
V-C Stream A: Forensic Psycholinguistics (CatBoost)
This branch processes tabular features derived from the text. We extracted 35 numerical features (Table I) and trained a CatBoost classifier [17].
  • Latency: $O(N\cdot d)$ where $d$ is the depth of the tree.
  • Interpretability: High (via SHAP values).
TABLE I: Psycholinguistic and Structural Feature Set
Category Count Examples
Psycholinguistic 12 Authority, Scarcity, Reciprocity
Forensic 7 Hedges, Boosters, Exclusive Words
Sentiment 6 Polarity, Subjectivity, $\psi$ Score
Structural 10 Caps Ratio, URL Count, Punctuation
V-D Stream B: Semantic Deep Learning (DistilBERT)
We use the non-cased DistilBERT-base [8] for contextual understanding. We performed the training on a Tesla T4 GPU (CUDA 12.6, PyTorch) with 14.7 GB VRAM allocation.
  • Complexity: $O(N^{2})$ due to the self-attention mechanism.
  • Input: Raw tokens (truncated to 512).
VI Experimental Setup
VI-A Dataset Composition
Table II details the balanced dataset (4,003 Legitimate / 3,987 Fraud). The test set is a stratified 20% hold-out.
TABLE II: Dataset Distribution (Balanced Enron + Synthetic BEC)
Subset Legitimate Fraudulent Total
Training (80%) 3,202 3,202 6,404
Testing (20%) 801 785 1,586
Total 4,003 3,987 7,990
VI-B Modern Data Validation (PMCC-2025)
To test temporal generalization, we curated a private corpus of 1,000 legitimate business emails from 2025 (PMCC-2025), containing modern slang and mobile signatures.
VI-C Hyperparameter Optimization
Table III lists the optimal parameters for CatBoost found via Optuna.
TABLE III: Optimized CatBoost Hyperparameters (Optuna)
Parameter Value
Iterations 545
Tree Depth 8
Learning Rate 0.0780
L2 Regularization 1.29
Bagging Temperature 3.10
VII Results and Analysis
TABLE IV: Comparative Performance on Tesla T4 GPU
Model Detection Metrics Efficiency
AUC F1 Recall Precision Latency (avg) Latency (P95)
CatBoost (Forensic) 0.9905 0.9486 0.9516 0.9456 0.885 ms 0.978 ms
DistilBERT (Semantic) 1.0000 0.9981 1.0000 0.9963 7.403 ms 10.873 ms
Speedup Factor 8.4x
VII-A Comparative Performance
Table IV presents the head-to-head comparison on Tesla T4 GPU. DistilBERT achieves perfect detection (AUC 1.0000, F1 0.9981) with 7.403 milliseconds inference latency per email. CatBoost achieves strong detection (AUC 0.9905, F1 0.9486) in 0.885 milliseconds per sample, closing the latency gap. DistilBERT now operates within real-time constraints acceptable for most Mail Transfer Agent implementations. CatBoost latency remains advantageous for ultra-high-throughput environments (exceeding 100,000 emails per hour) or CPU-only infrastructure.
Figure 2 shows the latency distribution across 1,586 test samples. CatBoost exhibits consistent sub-millisecond performance with minimal variance, while DistilBERT shows acceptable latency with some outliers due to variable input lengths.
Figure 2: Latency Distribution (1,586 samples). CatBoost demonstrates consistent sub-millisecond performance (mean = 0.885 ms, green triangle). DistilBERT achieves acceptable real-time latency (mean = 7.403 ms, green triangle) with GPU acceleration, though with higher variance due to variable text lengths.
VII-B Learning Dynamics
Figure 3 shows the learning curve for CatBoost. The model reaches optimal AUC performance with approximately 3,000 training examples, with validation performance plateauing at 0.9983 AUC. The training curve remains at 1.0000 throughout, indicating the model has sufficient capacity without overfitting due to proper regularization.
Figure 3: Learning Curve (CatBoost). Validation AUC plateaus at approximately 3,000 training samples. The gap between training (red) and validation (blue) curves remains minimal, indicating good generalization. The shaded region represents 95% confidence intervals across 5-fold cross-validation.
VII-C Robustness & Statistical Significance
Table V details the degradation under adversarial attacks. McNemar’s test ($\chi^{2}=18.38,p<0.001$) confirms the models have statistically distinct error profiles.
TABLE V: Robustness Analysis: Degradation Under Attack
Model Clean Recall Poisoned Recall Drop (%)
DistilBERT (Semantic) 100.00% 99.55% 0.4%
CatBoost (Forensic) 98.93% 97.31% 1.6%
Figure 4: Adversarial Robustness Analysis and Error Distribution. Top: Recall on Clean vs. Poisoned data (Left) and Degradation magnitude (Right). Bottom: McNemar’s Contingency Table (Left) and Pie Chart of Agreement (Right). The statistical analysis indicates that we must manage the significant disagreement (1.6%).
Figure 4 visualizes the error distribution. DistilBERT demonstrates superior robustness to homoglyph attacks. The Forensic model captures 98.4% of threats under adversarial conditions at substantially lower computational cost.
VII-D Model Agreement Analysis
Figure 5 shows the correlation between CatBoost and DistilBERT probability predictions. The high correlation coefficient (r = 0.9756) indicates that both models generally agree on classification decisions, with most disagreements occurring in the low-probability (clearly legitimate) region. The cluster of points at (0, 1) and (1, 1) in DistilBERT’s predictions reflects its higher confidence on fraudulent samples.
Figure 5: Model Prediction Correlation (r = 0.9756). Both models show strong agreement on most samples. The majority of samples cluster near (0, 0) for legitimate emails and (1, 1) for fraudulent emails. DistilBERT shows more decisive predictions (binary clustering), while CatBoost exhibits more graduated probability distributions.
VII-E Explainability
Figure 6 shows the top features driving the Forensic model, ranked by their discriminative power. Figure 7 reveals how these features influence predictions directionally. ent_money_count and tech_threat_count are dominant predictors.
Figure 6: Feature Importance Ranking. Money entity detection (ent_money_count) provides the strongest discriminative signal, followed by technical threat keywords and financial terminology. Importance is measured by the mean decrease in accuracy when a feature is permuted.
Figure 7: SHAP Feature Importance. Money entities (ent_money_count) and threat keywords (tech_threat_count) dominate the Forensic decision logic. Blue regions indicate features that push predictions toward legitimate classification, while red regions indicate fraud signals. The wide distribution of certain features (e.g., complex_word_ratio) shows they have variable impacts depending on context.
Figure 8 illustrates the DistilBERT attention mechanism, showing high impact on financial keywords.
Figure 8: DistilBERT Token Impact. High attention weights assigned to financial terminology and transfer keywords.
VIII Forensic Failure Analysis
VIII-A Case Study: The “Johns-Deleon” Attack
We analyze the single False Negative (Index 484) observed in Table VI.
TABLE VI: Confusion Matrix (CatBoost, Threshold = 0.12)
Predicted
Legitimate Fraud
True Legitimate 764 37
Fraud 1 784
The attack email read:
Subject: Pmnt for Johns-Deleon
Body: Sent from my phone.
Forensic Findings:
  1. 1.
    Homoglyph Evasion: The character ‘e’ in “phone” was replaced with Cyrillic Small Letter e ($U+0435$).
  2. 2.
    Feature Starvation: The message contained only 5 words. This brevity limited the analysis of language patterns (see Figure 9).
Figure 9: Feature Starvation Analysis. The False Negative (marked star) falls into the ultra-short message zone (fewer than 15 words), where linguistic signal is too weak.
VIII-B Decision Policy Algorithm
Based on this analysis, we propose the Zero-Trust Grey Zone Policy (Algorithm 2).
Algorithm 2 Zero-Trust Grey Zone Policy
0: Email $x$, Thresholds $\tau_{L},\tau_{H}$, Model $M$
1: Prob $p\leftarrow M(x)$
2: Length $L\leftarrow\text{count\_words}(x)$
3:if $L<15$ AND $x$ contains Financial Keyword then
4:  return Manual Review (Safeguard)
5:end if
6:if $p<\tau_{L}$ then
7:  return Auto-Allow
8:else if $p>\tau_{H}$ then
9:  return Auto-Block
10:else
11:  return Manual Review (Grey Zone)
12:end if
IX Economic Optimization
IX-A Calibration
Figure 10 shows the reliability diagram. The low Brier Score (0.0016) validates the probabilities for financial decision-making.
Figure 10: Reliability Diagram. The model is highly calibrated (Brier Score = 0.0016).
IX-B Cost Surface & ROI
Figure 11 illustrates the financial optimization landscape. The optimal policy is $\tau_{L}\approx 0$, maximizing threat capture while minimizing false positives.
Figure 11: Financial Optimization Surface. The “Valley of Safety” (dark blue region) minimizes total financial loss across threshold combinations.
As shown in Figure 12, the sensitivity analysis reveals that the return on investment remains substantial even when considering different fraud cost assumptions. This builds on our previous discussion of the cost-benefit analysis.
Figure 12: ROI Sensitivity Analysis. Defense effectiveness remains high across different fraud cost estimates.
IX-C GPU vs. CPU Infrastructure Trade-offs
Hardware configuration specifically influences inference latency on Tesla T4 GPU infrastructure:
  • DistilBERT: 7.403 milliseconds per email (batch size 1)
  • CatBoost: 0.885 milliseconds per email
For organizations processing 100,000 emails per hour:
  • DistilBERT infrastructure: approximately 205 GPU-hours per year (annual cost: \$150–300 on cloud platforms)
  • CatBoost infrastructure: approximately 25 CPU-hours per year (negligible cost)
Organizations should choose DistilBERT for GPU workloads or CatBoost for filtering.
X Conclusion
We evaluated the effectiveness of Semantic Deep Learning (DistilBERT) versus Forensic Psycholinguistics (CatBoost) for Business Email Compromise detection, specifically focusing on detection accuracy, inference latency, and resource efficiency on GPU-accelerated infrastructure (Tesla T4).
DistilBERT achieves superior detection (AUC 1.0000, F1 0.9981) with acceptable real-time latency (7.403 milliseconds) when GPU resources are available. CatBoost achieves competitive detection (AUC 0.9905, F1 0.9486) at 8.4x lower latency (0.885 milliseconds), consuming negligible GPU or CPU resources. For organizations with GPU infrastructure (cloud providers, large enterprises), DistilBERT provides maximum accuracy with acceptable latency, while for edge deployments, CPU-only gateways, or cost-sensitive environments, CatBoost delivers near-equivalent security with reduced operational overhead. This work demonstrates that the architecture choice depends on infrastructure availability, as semantic models with GPU acceleration provide maximum accuracy while forensic models are suited for resource-constrained scenarios. Both approaches exceed 99.96% accuracy.
Acknowledgment
The author thanks Dr. Frederick Ayivor, Ph.D, for guidance and support throughout this work. His advice on research design, analysis and interpretation helped shape the direction of the study. The author also thanks the open-source Kaggle community for providing datasets and tools that enabled data processing and initial experiments.
Appendix A Feature Definitions
TABLE VII: Psycholinguistic Feature Definitions
Feature Description
$\psi$ Score Interaction of politeness and urgency.
Authority Count of organizational titles and imperative cues.
Hedges Frequency of tentative expressions.
Sentiment $\Delta$ Shift in sentiment polarity across the email.
Caps Ratio Proportion of uppercase characters.

### v2

@@
 Abstract
-Business Email Compromise (BEC) is a sophisticated social engineering threat that manipulates organizational hierarchies and exploits psychological vulnerabilities like trust and authority, leading to significant financial and reputational damage for enterprises.
-According to the 2024 FBI Internet Crime Report, BEC accounts for over \$2.9 billion in annual adjusted losses, presenting a significant economic asymmetry: the cost of a False Negative (fraud loss) exceeds the cost of a False Positive (manual review) by orders of magnitude (approximately 1 to 5,480).
-This paper examines two distinct detection paradigms for BEC: the Forensic Psycholinguistic Stream, which utilizes CatBoost to analyze psycholinguistic cues with high interpretability and low latency, and the Semantic Stream, which employs DistilBERT for deep learning-based contextual language understanding, offering superior accuracy at the cost of higher computational resources.
-We evaluated DistilBERT on an adversarially poisoned dataset (N = 7,990) generated via our Black Hole protocol.
-It was benchmarked on Tesla T4 GPU infrastructure, achieving superior detection (AUC = 1.0000, F1 = 0.9981) with acceptable real-time latency (7.403 milliseconds).
-CatBoost achieves competitive detection (AUC = 0.9905, F1 = 0.9486) at 8.4x lower latency (0.885 milliseconds), consuming negligible computational resources.
-For organizations with GPU infrastructure, DistilBERT offers superior accuracy, making it an ideal choice.
-On the other hand, CatBoost is preferable for edge deployments or cost-sensitive environments due to its comparable security and lower operational costs, providing a viable alternative.
-Both approaches demonstrate a return on investment exceeding 99.96% when optimized through cost-sensitive learning, as evidenced by their ability to significantly reduce false negatives and associated financial losses.
-This high return is achieved by minimizing the financial impact of undetected fraudulent emails, which are substantially more costly than false positives.
+Business Email Compromise (BEC) is a sophisticated social engineering threat that manipulates organizational hierarchies, leading to significant financial damage.
+According to the 2024 FBI Internet Crime Report, BEC accounts for over \$2.9 billion in annual losses, presenting a massive economic asymmetry: the financial cost of a False Negative (fraud loss) exceeds the operational cost of a False Positive (manual review) by a ratio of approximately 5,480:1.
+This paper contrasts two detection paradigms: a Forensic Psycholinguistic Stream (CatBoost), which analyzes linguistic cues like urgency and authority with high interpretability, and a Semantic Stream (DistilBERT), which utilizes deep learning for contextual understanding.
+We evaluated both streams on a hybrid dataset (N=7,990) containing human-legitimate and AI-synthesized adversarial fraud.
+Benchmarked on Tesla T4 infrastructure, DistilBERT achieved near-perfect detection on synthetic threats (AUC >> 0.99, F1 = 0.998) with acceptable real-time latency (7.4 ms).
+CatBoost achieved competitive detection (AUC = 0.991, F1 = 0.949) at 8.4x lower latency (0.8 ms) with negligible resource consumption.
+We conclude that while DistilBERT offers maximum accuracy for GPU-equipped organizations, CatBoost provides a viable, cost-effective alternative for edge deployments.
+Both approaches demonstrate a theoretical ROI exceeding 99.9% when optimized via cost-sensitive learning.
 I Introduction
@@
   1. 1.
-    Economic Asymmetry: The cost of a False Positive ($C_{FP}$) is approximately \$25.
-The cost of a False Negative ($C_{FN}$) is approximately \$137,000 on average [1].
+    Economic Asymmetry: The operational reality is defined by cost, not just frequency.
+The cost of a False Positive ($C_{FP}$) is approximately \$25, while the cost of a False Negative ($C_{FN}$) is approximately \$137,000 on average [1].
 Standard loss functions fail to capture this 1:5,480 risk ratio.
@@
   • T1036.005 (Masquerading): Using homoglyphs to mimic legitimate keywords (e.g., “Payment”).
+III-C Limitations & Threats to Validity
+Synthetic fraud constitutes 52.7% of the dataset.
+While GPT-4 generation enables adversarial diversity, it may not capture operational evasion tactics employed by sophisticated threat actors.
+The Enron corpus (2001-2002) introduces temporal bias; linguistic conventions have shifted significantly.
+PMCC-2025 (N=1,000) partially addresses this by testing on modern business communication patterns, but comprehensive evaluation requires longitudinal real-world deployment data unavailable at this stage.
 IV Theoretical Framework
@@
 BEC attacks violate this while adhering to the Maxim of Manner (be polite) to mask the deception.
-We quantify this “Polite Coercion” using the Smiling Assassin Score ($\psi$).
+We operationalize this dynamic as an engineered feature—the Smiling Assassin Score ($\psi$)—capturing politeness-urgency interaction.
+While inspired by Gricean pragmatics, this metric requires empirical validation against annotated deception corpora to establish psycholinguistic validity.
 Let $S_{pos}(x)\in[0,1]$ be the normalized positive sentiment and $U_{\text{freq}}(x)$ be the urgency token density.
@@
 To test temporal generalization, we curated a private corpus of 1,000 legitimate business emails from 2025 (PMCC-2025), containing modern slang and mobile signatures.
+PMCC-2025 contains contemporary linguistic patterns absent from the Enron-based test set: mobile signatures, emoji usage, informal syntax.
+This validates temporal generalization models trained on 2001-2002 data must handle 2025 communication styles without catastrophic drift.
 VI-C Hyperparameter Optimization
@@
 Figure 12: ROI Sensitivity Analysis. Defense effectiveness remains high across different fraud cost estimates.
+The optimal threshold ($\tau_{L}\approx 0$) maximizes financial utility under fixed cost assumptions.
+However, operational deployment introduces analyst workload constraints and alert fatigue risks not captured in this model.
+Organizations processing 100K+ emails daily may require higher thresholds to maintain sustainable review queues, trading marginal detection gains for operational feasibility.
 IX-C GPU vs. CPU Infrastructure Trade-offs
@@
 Both approaches exceed 99.96% accuracy.
+Future work must validate performance on longitudinal operational data to address synthetic bias and temporal drift.
+Psycholinguistic feature constructs require empirical validation through controlled deception studies.
 Acknowledgment
-The author thanks Dr.
-Frederick Ayivor, Ph.D, for guidance and support throughout this work.
-His advice on research design, analysis and interpretation helped shape the direction of the study.
-The author also thanks the open-source Kaggle community for providing datasets and tools that enabled data processing and initial experiments.
+The authors thank the open-source Kaggle community for providing datasets and tools that enabled data processing and initial experiments.
 Appendix A Feature Definitions
+Table VII lists the psycholinguistic features used in the analysis.
+It defines each feature and states what linguistic attribute it captures, including politeness–urgency interaction, authority cues, hedge frequency, sentiment shift, and uppercase character ratio.
 TABLE VII: Psycholinguistic Feature Definitions